Calculate the energy of a photon with wavelength of 4.50 × 103 nm.
1
Expert's answer
2019-10-21T07:16:07-0400

E=h*c/"\lambda"

h=6.626 * 10^-34 Kg*m2/s

c=3*10^8 m/s

λ = 4.5*103*10^(-9)=463.5*10^(-9) m

E= 6.626*10^(-34) *3*10^(8)/463.5/10*(-9)=4.29*10^(-19) eV = 0.669 Joules' ( 1 ev = 6.409*10^(-19) J)